OpenStudy (anonymous):

A particular strain of bacteria doubles in population every 10 minutes. Assuming you start with a single bacterium in a petri dish, how many bacteria will there be in 2.5 hours? A. 9536 B. 32,768 C. 65,536 D. 16,384

OpenStudy (anonymous):

its b right???

OpenStudy (kash_thesmartguy):

2.5hrs=150 min. 2^14=D.

OpenStudy (anonymous):

so i was wrong then T_T

OpenStudy (anonymous):

I think you were right. 2^15

OpenStudy (kash_thesmartguy):

how does 2^15 because it starts from 1 right?

OpenStudy (anonymous):

0 min = 1 = 2^0 10 min = 2^1 20 min = 2^2 30 min = 2^3 40 min = 2^4 50 min = 2^5 60 min = 2^6 70 min = 2^7 80 min = 2^8 90 min = 2^9 100 min = 2^10 110 min = 2^11 120 min = 2^12 130 min = 2^13 140 min = 2^14 150 min = 2^15

OpenStudy (anonymous):

so i was right

OpenStudy (kash_thesmartguy):

yeah, that's right^

OpenStudy (anonymous):

it appears so

OpenStudy (kash_thesmartguy):

Sry :(

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

@Kash_TheSmartGuy right process, just miscounted